maximum height reached by projectile is 4m. The horizontal range is 12m. velocity of projection?

Grade:12

1 Answers

Arun
6 years ago
Dear Student
 
H  = (R tan Ф ) / 4 
4  = (12 tan Ф ) / 4

tan Ф = 4/3        sin 2Ф = 2 tan Ф / (1 + tan² Ф)  = 0.96

R = u² sin 2Ф / g  

u² =  12 * g / 0.96  = 25 g /2 

u = 5 √(g/2)
 
Regards
Arun (askIITians forum expert)